OpenStudy (anonymous):

4x^3/3x * 14/3 x^2-1/x divided by x+1/x-1

OpenStudy (anonymous):

help?!?!

OpenStudy (anonymous):

Please, use brackets and pharenthesis so that we can understand what the prob we need to do..lol

OpenStudy (anonymous):

seriously, you don't get it? \[\left(\begin{matrix}x ^{2} -1\\ x\end{matrix}\right)\div \left(\begin{matrix}x+1 \\ x-1\end{matrix}\right)\]

OpenStudy (anonymous):

is this like this? (x^2-1)/x divided by( x+1)/(x-1) ? or like this (x^2)-(1/x) divided by (x+1)/(x-1) i will assume it like this, if (x^2-1)/x divided by( x+1)/(x-1) (x^2-1) (x-1) (x-1)(x+1) x-1 =------- * -------- = --------- * ----------- cancel the x+1 x ( x+1) x x+1 (x-1)(x-1) x^2-2x+1 = -------------= ------------ = x-2+(1/x) ans x x

OpenStudy (anonymous):

would that help you?

OpenStudy (anonymous):

not sure trying to type it out to see if it looks right

OpenStudy (anonymous):

is that the final answer?